How much water must be added to 30 litres of milk at 2 litres for $1.40 so as to have a mixture worth $0.50 a litre ?
Answer by Theo(13342)

30 liters of milk at 2 liters for $1.40 is worth 30 * $1.40 / 2 = $21.00

your cost per liter is equal to $21.00 / 30 = $.70 per liter.

You want your mixture to be worth $.50 per liter.

let x = the number of liters of water you need to add.

you want .50 * (x + 30) = $21.00

that's x + 30 liters of milk and water at $.50 per liter is equal to $21.00 which is the same value as 30 liters of pure milk milk at $.70 per liter.

your formula is:

$.50 * (x + 30) = $21.00

simplify to get:

$.50 * x + $15.00 = $21.00

subtract $15.00 from both sides of this equation to get:

$.50 * x = $6.00

divide both sides of this equation by $.50 to get:

x = 12

you need to add 12 liters of water to get a mixture with a value of $21.00.

total liters required = 30 milk + 12 water = 42 liters of mixture.

$.50 * 42 = $21.00 for 42 liters of milk and water mixture.